OpenStudy (anonymous):

i will fan and medal!! Circle A has a diameter of 8 inches, a circumference of 25.12 inches, and an area of 50.24 square inches. The diameter of circle B is 3 inches, the circumference is 9.42 inches, and the area is 7.065 square inches. Part A: Using the formula for circumference, solve for the value of pi for each circle. (4 points) Part B: Use the formula for area and solve for the value of pi for each circle. (4 points) Part C: What observation can you make about the value of pi for circles A and B? (2 points)

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Legends

OpenStudy (anonymous):

@DarkMoonZ

OpenStudy (legends):

@jigglypuff314

OpenStudy (michele_laino):

please the circumference C is given by the subsequent formula: \[C = \pi D\] where D is the diameter furthermore the area A of a circle is given by the subsequent formula: \[A = \pi \frac{{{D^2}}}{4}\] where D is the diameter of the circle

jigglypuff314 (jigglypuff314):

Circumference = (pi)*(diameter) they gave you Circumference of Circle A = 25.12 and Diameter of Circle A = 8 so (pi) = (Circumference) divided by (Diameter) = 25.12 / 8 = ?

OpenStudy (anonymous):

i have no idea what that means

OpenStudy (anonymous):

0.265?

jigglypuff314 (jigglypuff314):

try again? :) 25.12 divided by 8

OpenStudy (anonymous):

ooh 3.14

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

so part a is 3.14

jigglypuff314 (jigglypuff314):

mhmm I think it wants you to show the work?

OpenStudy (anonymous):

I will but i need help with part b and c too plz @jigglypuff314

jigglypuff314 (jigglypuff314):

Formula for Area of Circle = (pi) * (radius)^2 Radius = (1/2) * (Diameter) Diameter of Circle A = 8 so Radius = (1/2) * 8 = 4 Area of Circle A = 50.24 (pi) = (Area of Circle A) divided by (radius times radius)

OpenStudy (anonymous):

3.14

OpenStudy (anonymous):

so part b is 3.14 also?

jigglypuff314 (jigglypuff314):

yes

OpenStudy (anonymous):

Part C is? @nincompoop